# Yoga for Fourier–Mukai partnership

### v1

@@ 6. Applications
In this section, we study the consequences of utilizing Theorems˜4.6 and 5.4 together.
@@
Let $\overline{k}$ be the algebraic closure of $k$. Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{k}Y_{2})$ has an associated integral transform $\Phi_{K}$ restricting to an equivalence $D^{b}_{\operatorname{coh}}(C)\to D^{b}_{\operatorname{coh}}(C^{\prime})$. Denote by $\pi\colon(Y_{1}\times_{k}Y_{2})\times_{k}\operatorname{Spec}(\overline{k})\to Y_{1}\times_{k}Y_{2}$ the natural morphism. From Theorem˜4.6, we see that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $D^{b}_{\operatorname{coh}}(C\times_{k}\operatorname{Spec}(\overline{k}))\to D^{b}_{\operatorname{coh}}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$. Then $C\times_{k}\operatorname{Spec}(\overline{k})$ and $C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$ are Fourier–Mukai partners over an algebraically closed field of characteristic zero. By [CNS25, Corollary 5.4], we have that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $\operatorname{Perf}(C\times_{k}\operatorname{Spec}(\overline{k}))\to\operatorname{Perf}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$. From the following facts, we know that $C\times_{k}\operatorname{Spec}(\overline{k})\cong C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$:
@@
  • $C\times_{k}\operatorname{Spec}(\overline{k})$ Gorenstein of arithmetic genus one and trivial canonical line bundle [LM14, Theorem 1.1] ($C\times_{k}\operatorname{Spec}(\overline{k})$ is integral with arithmetic genus one such that $\operatorname{deg}(\omega_{X})=0$ implies $\omega_{X}\cong\mathcal{O}_{X}$ by [Cat82, Proposition 1.11] where $\omega_{X}$ is the canonical sheaf).
Observe these are the only possible cases for projective curves over an algebraically closed field in characteristic zero. Hence, there is a finite field extension $L/k$ such that $C\times_{k}\operatorname{Spec}(L)\cong C^{\prime}\times_{k}\operatorname{Spec}(L)$ (see e.g. [Poo17, Proposition 3.2.5]), which completes the proof. ∎
Example 6.2.
There is a pair of non-isomorphic elliptic curves $E_{1}$ and $E_{2}$ over $\mathbb{Q}$ that are Fourier–Mukai partners (see [War14, Example 2.3.11]). From [Orl02, Theorem 2.2], it follows that any triangulated equivalence $D^{b}_{\operatorname{coh}}(E_{1})\to D^{b}_{\operatorname{coh}}(E_{2})$ is given by an integral transform. Note that an elliptic curve is geometrically integral and projective. By Proposition˜6.1, there is a finite field extension $L/\mathbb{Q}$ such that $E_{1}\times_{\mathbb{Q}}\operatorname{Spec}(L)\cong E_{2}\times_{\mathbb{Q}}\operatorname{Spec}(L)$.
@@
Let $S$ be a separated Noetherian $\mathbb{Q}$-scheme. Suppose $f_{i}\colon Y_{i}\to S$ are proper flat morphisms whose special fibers are geometrically integral curves. Assume $f_{1}$ is locally projective. If $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$, then then genii of the special fibers of $Y_{1}$ coincide with that of $Y_{2}$.
@@
implies that $\mathbb{Q}\subseteq\kappa(s)$ (i.e. $\kappa(s)$ has characteristic zero). Note that, as $S$ is separated, $\operatorname{Spec}(\kappa(s))\to S$ is affine (see e.g. [Sta25, Tag 01SG]). Denote by $Y_{i}^{s}$ for the base change of $Y_{i}$ along $\operatorname{Spec}(\kappa(s))\to S$. Our hypothesis ensures that each $Y_{i}^{s}$ is a geometrically integral proper curve over $\kappa(s)$. Then, by Theorem˜5.4, $Y^{s}_{1}$ and $Y_{2}^{s}$ are Fourier–Mukai partners over $\kappa(s)$. Furthermore, Proposition˜6.1 guarantees that there is a finite field extension $K(s)/\kappa(s)$ such that $Y^{s}_{1}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ is isomorphism to $Y^{s}_{2}\times_{\kappa(s)}\operatorname{Spec}(K(s))$. Hence, if coupled with [Sta25, Tag 0BY9 & Tag 0FD2], the genii of each $Y^{s}_{i}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ coincides with $Y^{s}_{j}$ for $i\not=j$. In other words, the genii of $Y^{s}_{1}$ and $Y_{2}^{s}$ coincide, which completes the proof. ∎
@@
Let $Y_{1}$ and $Y_{2}$ be proper schemes over a field $k$ (which need not be perfect). Suppose $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners. If $Y_{1}$ is geometrically regular, then so is $Y_{2}$. In other words, smoothness is a derived invariance.
@@
Let $f_{i}\colon Y_{i}\to S$ be locally projective flat morphisms to a separated Noetherian scheme. Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$. Then $f_{1}$ is smooth if, and only if, $f_{2}$ is smooth.
@@
We prove only one direction as the other follows nearly verbatim. Assume $f_{1}$ is smooth. It suffices, by [Sta25, Tag 01V8], to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are smooth over $\kappa(s)$ where $s\in S$ is any point. Fix $s\in S$ (perhaps, not closed). The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g. [Sta25, Tag 01SG]) and of finite tor-dimension (due to being flat). Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners. Then Theorem˜5.4 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$. However, smoothness is stable under base change, so each base change of $Y_{1}$ along such morphisms remains smooth. Hence, the desired claim follows from Proposition˜6.4. ∎
@@
Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms to a separated Noetherian scheme with geometrically integral fibers. Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$. Then $f_{1}$ is Cohen–Macaulay (resp. Gorenstein) if, and only if, $f_{2}$ satisfies the same condition.
@@
It suffices to show $f_{2}$ is Cohen–Macaulay (resp. Gorenstein) if $f_{1}$ satisfies the same condition because the proof is analogous for the converse. In fact, we argue in a similar fashion to Corollary˜6.5. Moreover, by [Sta25, Tag 045O] (resp. [Sta25, Tag 0C03]), we only need to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are Cohen–Macaulay (resp. Gorenstein) over an algebraic closure of each $\kappa(s)$ where $s\in S$ is any point. Fix $s\in S$ (perhaps, not closed). The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g. [Sta25, Tag 01SG]) and of finite tor-dimension (due to being flat). Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners. Then Theorem˜5.4 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$. Furthermore, if we base change further along $\overline{\kappa(s)}/\kappa(s)$ with $\overline{\kappa(s)}$ an algebraic closure of $\kappa(s)$, these base changes remain Fourier–Mukai partners via Theorem˜4.6. However, it follows (as fibers are geometrically integral) from [HLS09, Theorem 4.4], that these integral projective schemes over $\overline{\kappa(s)}$ enjoy the property that Cohen–Macaulayness (resp. Gorensteinnes) is a derived invariance. This completes the proof. ∎

### v2

@@ 5. Applications
In this section, we study the consequences of Theorem˜4.6. By ‘genus’ of a curve, we mean ‘arithmetic’(see e.g. [Sta25, Tag 0BY6]).
@@
Let $\overline{k}$ be the algebraic closure of $k$. Suppose $K\in D^{b}_{\operatorname{coh}}(C\times_{k}C^{\prime})$ has an associated integral transform $\Phi_{K}$ restricting to an equivalence $D^{b}_{\operatorname{coh}}(C)\to D^{b}_{\operatorname{coh}}(C^{\prime})$. Denote by $\pi\colon(C\times_{k}C^{\prime})\times_{k}\operatorname{Spec}(\overline{k})\to C\times_{k}C^{\prime}$ the natural morphism. From Theorem˜4.6, we see that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $D^{b}_{\operatorname{coh}}(C\times_{k}\operatorname{Spec}(\overline{k}))\to D^{b}_{\operatorname{coh}}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$. Then $C\times_{k}\operatorname{Spec}(\overline{k})$ and $C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$ are Fourier–Mukai partners over an algebraically closed field of characteristic zero. By [CNS25, Corollary 5.4], we have that $\Phi_{\mathbf{L}\pi^{\ast}K}$ restricts to an equivalence $\operatorname{Perf}(C\times_{k}\operatorname{Spec}(\overline{k}))\to\operatorname{Perf}(C^{\prime}\times_{k}\operatorname{Spec}(\overline{k}))$. Consequently, we know that $C\times_{k}\operatorname{Spec}(\overline{k})\cong C^{\prime}\times_{k}\operatorname{Spec}(\overline{k})$ in the following cases:
@@
  • $C\times_{k}\operatorname{Spec}(\overline{k})$ Gorenstein of genus one and trivial canonical line bundle [LM14, Theorem 1.1] ($C\times_{k}\operatorname{Spec}(\overline{k})$ is integral with genus one such that $\operatorname{deg}(\omega_{X})=0$ implies $\omega_{X}\cong\mathcal{O}_{X}$ by [Cat82, Proposition 1.11] where $\omega_{X}$ is the canonical sheaf).
Observe these are the only possible cases for projective curves over an algebraically closed field in characteristic zero (see e.g. the two paragraphs above [Spe23, Theorem A] or use [Cat82, Theorem B, Proposition 1.8 and Proposition 1.11]). Hence, there is a finite field extension $L/k$ such that $C\times_{k}\operatorname{Spec}(L)\cong C^{\prime}\times_{k}\operatorname{Spec}(L)$ (see e.g. [Poo17, Proposition 3.2.5]), which completes the proof. ∎
@@
Let $S$ be a quasi-compact separated regular $\mathbb{Q}$-scheme. Suppose $f_{i}\colon Y_{i}\to S$ are proper flat morphisms whose special fibers are geometrically integral curves. Assume $f_{1}$ is locally projective. If $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$, then then genii of the special fibers of $Y_{1}$ coincide with that of $Y_{2}$.
@@
implies that $\mathbb{Q}\subseteq\kappa(s)$ (i.e. $\kappa(s)$ has characteristic zero). Note that, as $S$ is separated, $\operatorname{Spec}(\kappa(s))\to S$ is affine (see e.g. [Sta25, Tag 01SG]). Denote by $Y_{i}^{s}$ for the base change of $Y_{i}$ along $\operatorname{Spec}(\kappa(s))\to S$. Our hypothesis ensures that each $Y_{i}^{s}$ is a geometrically integral proper curve over $\kappa(s)$. Then, by Theorem˜4.6, $Y^{s}_{1}$ and $Y_{2}^{s}$ are Fourier–Mukai partners over $\kappa(s)$. Indeed, the hypothesis on $S$ ensures it is Noetherian, and so the closed immersion $\operatorname{Spec}(\kappa(s))\to S$ is locally of finite tor-dimension (see e.g. [GW23, Remark 25.39(2)]). Furthermore, Proposition˜5.1 guarantees that there is a finite field extension $K(s)/\kappa(s)$ such that $Y^{s}_{1}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ is isomorphism to $Y^{s}_{2}\times_{\kappa(s)}\operatorname{Spec}(K(s))$. Hence, if coupled with [Sta25, Tag 0BY9 & Tag 0FD2], the genii of each $Y^{s}_{i}\times_{\kappa(s)}\operatorname{Spec}(K(s))$ coincides with $Y^{s}_{j}$ for $i\not=j$. In other words, the genii of $Y^{s}_{1}$ and $Y_{2}^{s}$ coincide, which completes the proof. ∎
@@
Let $Y_{1}$ and $Y_{2}$ be proper schemes over a field $k$ (which need not be perfect). Suppose $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners. If $Y_{1}$ is geometrically regular, then so is $Y_{2}$. In other words, smoothness is an invariance for Fourier-Mukai partners.
@@
Let $f_{i}\colon Y_{i}\to S$ be locally projective flat morphisms to a quasi-compact separated regular scheme. Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$. Then $f_{1}$ is smooth if, and only if, $f_{2}$ is smooth.
@@
We prove only one direction as the other follows nearly verbatim. Assume $f_{1}$ is smooth. It suffices, by [Sta25, Tag 01V8], to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are smooth over $\kappa(s)$ where $s\in S$ is any point. Fix a point $s\in S$ (which need not be closed). The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g. [Sta25, Tag 01SG]) and of finite tor-dimension (due to being flat). Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners. A further application Theorem˜4.6 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$. Indeed, $S$ being regular implies $\mathcal{O}_{S,s}$ is as well, and so the closed immersion from $\operatorname{Spec(\kappa(p))}$ is of finite tor-dimension (see e.g. [GW23, Remark 25.39(2)]). However, smoothness is stable under base change, so each base change of $Y_{1}$ along such morphisms remains smooth. Hence, the desired claim follows from Proposition˜5.3. ∎
@@
Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms with geometrically integral fibers. Suppose $S$ is quasi-compact, separated, and regular. Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$. Then $f_{1}$ is Cohen–Macaulay (resp. Gorenstein) if, and only if, $f_{2}$ satisfies the same condition.
@@
It suffices to show $f_{2}$ is Cohen–Macaulay (resp. Gorenstein) if $f_{1}$ satisfies the same condition because the proof is analogous for the converse. In fact, we argue in a similar fashion to Corollary˜5.4. Moreover, by [Sta25, Tag 045O] (resp. [Sta25, Tag 0C03]), we only need to show that all the fibers $Y_{2}\times_{S}\operatorname{Spec}(\kappa(s))$ are Cohen–Macaulay (resp. Gorenstein) over an algebraic closure of each $\kappa(s)$ where $s\in S$ is any point. Fix $s\in S$ (perhaps, not closed). The natural morphism $t_{s}\colon\operatorname{Spec}(\mathcal{O}_{S,s})\to S$ is affine (see e.g. [Sta25, Tag 01SG]) and of finite tor-dimension (due to being flat). Hence, Theorem˜4.6 tells us the base changes of $Y_{1}$ and $Y_{2}$ along $t_{s}$ are Fourier–Mukai partners. Then Theorem˜4.6 implies these base changes remain Fourier–Mukai partners if we base change further along the natural closed immersion $\operatorname{Spec}(\kappa(p))\to\operatorname{Spec}(\mathcal{O}_{S,s})$. Furthermore, if we base change further along $\overline{\kappa(s)}/\kappa(s)$ with $\overline{\kappa(s)}$ an algebraic closure of $\kappa(s)$, these base changes remain Fourier–Mukai partners via Theorem˜4.6. However, it follows (as fibers are geometrically integral) from [HLS09, Theorem 4.4], that these integral projective schemes over $\overline{\kappa(s)}$ enjoy the property that Cohen–Macaulayness (resp. Gorensteinnes) is a derived invariance. This completes the proof. ∎
Under additional conditions, we can improve Proposition˜5.5 in the Gorenstein case.
Proposition 5.6.
Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms to a separated Noetherian scheme, where $Y_{1}$ and $Y_{2}$ are of finite Krull dimension, satisfying the resolution property (see e.g. [Sta25, Tag 0F86]). Assume that $Y_{1}$ and $Y_{2}$ are derived equivalences (e.g. the equivalence need not arise from an integral transform). Then $f_{1}$ is Gorenstein if, and only if, $f_{2}$ satisfies the same condition.
Proof.
We prove the case where $f_{1}$ is Gorenstein implies that of $f_{2}$ because the proof for the converse holds verabtim. Denote by $D^{b}_{\operatorname{fid}}(Y_{2})$ the full triangulated subcategory in $D^{b}_{\operatorname{coh}}(Y_{2})$ of complexes with bounded cohomology and with a bounded injective resolution (i.e. can be represented by a complex with finitely many nonzero components where each is injective $\mathcal{O}_{X}$-module). By [Har66, Subchapter V.2], the dualizing sheaf $f_{2}^{!}\mathcal{O}_{S}$ is contained in $D^{b}_{\operatorname{fid}}(Y_{2})$. Furthermore, [Sta25, Tag 0A89] tells us that there is an equivalence
(5.1) $D:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(-,f_{2}^{!}\mathcal{O}_{S})\colon D^{b}_{\operatorname{coh}}(Y_{2})\xrightarrow{\sim}D^{b}_{\operatorname{coh}}(Y_{2})^{\operatorname{op}}$
satisfying $D\circ D=\operatorname{id}$. Moreover, this equivalence restricts to an equivalence [Har66, Proposition V.2.6 a)],
(5.2) $D\colon\operatorname{Perf}(Y_{2})\xrightarrow{\sim}D^{b}_{\operatorname{fid}}(Y_{2})^{\operatorname{op}}$
So, we have the following intrinsic characterization of $D^{b}_{\operatorname{fid}}(Y_{1})$ in $D^{b}_{\operatorname{coh}}(Y_{1})$,
(5.3) $D^{b}_{\operatorname{fid}}(Y_{1})=\{B\in D^{b}_{\operatorname{coh}}(Y_{1})\mid\operatorname{Hom}(A,B[p])=0,\text{for}\ 0\ll|p|\text{ and all }A\in D^{b}_{\operatorname{coh}}(Y_{1})\}.$
Indeed, by [Orl06, Proposition 1.11], we have a dual characterization for perfect complexes in $D^{b}_{\operatorname{coh}}(Y_{1})$,
$\operatorname{Perf}(Y_{1})=\{B\in D^{b}_{\operatorname{coh}}(Y_{1})\mid\operatorname{Hom}(B,A[p])=0,\text{for}\ 0\ll|p|\text{ and all }A\in D^{b}_{\operatorname{coh}}(Y_{1})\},$
because $Y_{1}$ is separated Noetherian of finite Krull dimension and satisfies the resolution property (Orlov calls this (ELF) scheme - note that the proof in loc. cit. does not use that the scheme is defined over a field). This characterization of $\operatorname{Perf}$, together with the equivalences (5.1) and (5.2), implies (5.3).
Recall that $Y_{1}$ over $S$ is Gorenstein if, and only if, the dualizing sheaf $f_{1}^{!}\mathcal{O}_{S}$ is invertible (see e.g. [Har66, Exercise V.9.7]). In particular, $f_{1}^{!}\mathcal{O}$ is a perfect complex. If now $\Phi\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ is an equivalence, then $\Phi(f_{1}^{!}\mathcal{O})$ is a perfect complex in $D_{\operatorname{qc}}(Y_{2})$. Furthermore, as $\Phi$ restrict to an equivalence on $D^{b}_{\operatorname{coh}}$, we see by the characterization (5.3) that $\Phi(f_{1}^{!}\mathcal{O})$ is contained in $D^{b}_{\operatorname{fid}}(Y_{2})$. Hence, by (5.2), we see that $\Phi(f_{1}^{!}\mathcal{O})$ is of the form
$\Phi(f_{1}^{!}\mathcal{O})\cong D(P)=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(P,f_{2}^{!}\mathcal{O})\cong P^{\vee}\otimes^{\mathbf{L}}f_{2}^{!}\mathcal{O}\in D^{b}_{\operatorname{fid}}(Y_{2})$
for some $P$ in $\operatorname{Perf}(Y_{2})$. However, $\Phi(f_{1}^{!}\mathcal{O})$, and so, $P^{\vee}\otimes^{\mathbf{L}}f_{2}^{!}\mathcal{O}$ is in $\operatorname{Perf}(Y_{2})$. Hence, it follows that $f_{2}^{!}\mathcal{O}$ is also in $\operatorname{Perf}(Y_{2})$. Moreover, since $f_{2}^{!}\mathcal{O}^{\vee}\otimes^{\mathbf{L}}f_{2}^{!}\mathcal{O}\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(f_{2}^{!}\mathcal{O},f_{2}^{!}\mathcal{O})\cong\mathcal{O}_{Y_{2}}$ by the definition of a dualizing complex, it is easy to deduce that $f_{2}^{!}\mathcal{O}$ is a shift of a line bundle of $Y_{2}$ (see e.g. [Har66, Lemma V.3.3]). Again, by [Har66, Exercise V.9.7], we conclude that $Y_{2}$ over $S$ is Gorenstein, which completes the proof. ∎